OpenStudy (faman39):

Kenny goes shopping for school clothes during a tax−free weekend, and he saves 6% sales tax. Use a percent proportion to determine how much he would have to spend to save 30% . Which of the following is true? The solution is $18 , and is a denominator in the proportion. The solution is $18, and is a numerator in the proportion. The solution is $500, and is a denominator in the proportion. The solution is $500, and is a numerator in the proportion.

ganeshie8 (ganeshie8):

according to the question, we must setup a proportion like this : \(\frac{a}{b} = \frac{c}{d}\)

ganeshie8 (ganeshie8):

since he save 6 rupees on every 100 rupees he spends, first ratio will be 6/100 \(\frac{6}{100} = \frac{c}{d}\)

ganeshie8 (ganeshie8):

call : x, the amount he will have to spend

ganeshie8 (ganeshie8):

to save 30 dollars, the proportion becomes : \(\frac{6}{100} = \frac{30}{x}\)

ganeshie8 (ganeshie8):

solving gives, x = 500, and it is in the numerator.....

ganeshie8 (ganeshie8):

and i see theres a mistake in the question : its not 30%, it should be 30 $.

OpenStudy (faman39):

Thank u so much @ganeshie8 !!!!!!!!!!!!!!!!!!!!!!!!!!! :)

ganeshie8 (ganeshie8):

yw... :))

OpenStudy (faman39):

Yeah, sorry it ws my mistake :)

OpenStudy (faman39):

i m sorry about that

ganeshie8 (ganeshie8):

lol okay... its easy to make out so no probs... :D

OpenStudy (faman39):

your awesome @ganeshie8 !!!!!!!!!! :)

OpenStudy (rsadhvika):

minor correction @ganeshie8 . x is 500 and it was in the denominator

ganeshie8 (ganeshie8):

oops! you're right.... good catch @rsadhvika thank you :)